OpenStudy (anonymous):

find the additive inverse (opposite) of 18/23

OpenStudy (theeric):

Additive inverse is just the number you have to add to get 0, right? Well, you can solve for it, but pretty soon you'll just be seeing that it's very simple to do. \[\frac{18}{23} + t_{additiveInverseNumberYouWant}=0\]

OpenStudy (anonymous):

23/18 = 1 and 5/18

OpenStudy (theeric):

That's the reciprocal, @countonme ! :)

OpenStudy (theeric):

@countonme123 *

OpenStudy (anonymous):

lol my bad

OpenStudy (anonymous):

it would be -18/23

OpenStudy (theeric):

No problem! So many words in math.. You wouldn't think it right away. But there are.

OpenStudy (theeric):

vikrantg4 is right.

OpenStudy (theeric):

18/23 + number = 0 So number = 0 - 18/23 = -18/23. And it's always true that the additive inverse of a number will always be it's negative. "additive inverse property" means that you can add the negative of a number, to that number, to get 0.

OpenStudy (theeric):

In case you want to see that it is ALWAYS true, here you go. x + number = 0 number = 0 - x = -x. That's it, simply. Or, x + (-x) = 0. See, x is opposite of -x - and that's ANY x; so it's proved.